sovle simultaneously for x and y
2y+3x=5 and 3x-3y=x+12

Answer by Cromlix(4381)

Start by rearranging equations
3x + 2y = 5 ....1
2x - 3y = 12....2
Multiply (1) by 3
and (2) by 2
9x + 6y = 15.....3
4x - 6y = 24.....4
Add (3) + (4)
13x = 39
x = 3
Substitute x = 3 in (1)
3(3) + 2y = 5
9 + 2y = 5
2y = 5 - 9
2y = -4
y = -2
x = 3: y = -2
Hope this helps
:-)